OpenStudy (lgbasallote):

A stone is thrown from the top of a building with an initial velocity of 20 m/s downward. The top of the building is 60m above the ground. How much time elapses between the instant of release and the instant of impact with the ground?

OpenStudy (lgbasallote):

i did \[\implies 60 = 20t^2 + 4.8t\]

OpenStudy (lgbasallote):

then my calulator said t = 1.6 but the right answer was t - 2

OpenStudy (lgbasallote):

t = 2*

OpenStudy (lgbasallote):

can anyone explain where i went wrong?

OpenStudy (ghazi):

9.8/2=4.9

OpenStudy (lgbasallote):

uhh oops

OpenStudy (lgbasallote):

still doesnt give right answer

OpenStudy (ghazi):

but i got it 2

OpenStudy (lgbasallote):

how?

hartnn (hartnn):

which equation u used?

OpenStudy (ghazi):

\[t=\frac{ -20 \pm \sqrt{(20)^2+4*4.9*60} }{ 9.8 }=2.01\]

OpenStudy (lgbasallote):

\[x = V_1 t^2 + \frac{at}2\]

OpenStudy (ghazi):

and the equation is \[4.9t^2+20t-60=0\]

OpenStudy (lgbasallote):

i think i got it backwards...

OpenStudy (lgbasallote):

hmm no wonder

OpenStudy (lgbasallote):

i did have it backwards

hartnn (hartnn):

u took a,b,c values incorrect @ghazi

OpenStudy (ghazi):

@hartnn it is correct

OpenStudy (anonymous):

use s= ut + 1/2 gt^2...

hartnn (hartnn):

okk, i thought lg's equation was correct.

OpenStudy (lgbasallote):

i interchanged my signs incorrectly..that's why i cant get 2